- 8/18/16: Ben and Ed replace dome rollers after more rotation failures
- 5/30/14: Today we lubed the dome wheels and and track with a 3in1 Garage Door Lube.
We installed a permanent grease fitting. It seems strange to us that there is just one grease port but this must be a small bearing. We greased the azimuth bearing with Mobil 460. Uncertain how much grease to add we rotated the telescope by hand in azimuth as we added 20 pumps with the grease gun. After 360 degrees of rotation, we added another 10 and then parked the telescope.
- 7/9/13: reinstalled all three mirrors after coating and overcoating
by ATOC in Albuquerque. All went fairly smoothly. Used print overlay
of secondary to put a spot in physical center of mirror.
Did alignment of secondary
with collimating telescope, only adjust tilt a little. Did alignment of tertiary
with collimating telescope: adjusted tertiary rotation, and piston a bit.
Did initial primary aligment on sky. Note azimuth scale change because of
new az encoder shaft.
- 6/13: Ben rebuilt all azimuth idler boxes. Readjusted azimuth motor.
Replaced aluminum encoder cam with steel one.
- 8/24/11: dewar removed and put on pump, appears to have been very
soft. It has been warm most of the summer because of the filter wheel
problem and the weather.
- 8/24/11: Ed/Ben repair filter wheel which has been failing since
late June. Ed: "1-m Filter wheel detent problem is fixed. The original
symptoms were that the 1m would not rotate the filter wheel because
the 'DETENT OUT' signal was not being received by the OMS controller.
The detent motor could be heard turning so attention focused on the
'DETENT OUT' limit switch. I was able to measure switch action at the
detent connector by using a very long probe to actuate the switches.
Both checked out ok for switch function. This led me to troubleshoot the
system and interconnections. Had some problems due to poor documentation
but broke through when Ben built a breakout cable for the OMS (Oregon
Motion Systems) and we were able to stimulate the status bits manually
and confirm that the OMS side was working correctly. We verified the
cabling so that left the limit switch itself (even though the very
first test did not agree). Ben and Bill removed the unit and the bench
inspection revealed that the switches were indeed functioning well.
We reverse documented the circuit and saw that the 'DETENT OUT' switch
was not actuating when the detent was indeed out. Ben discovered that
the small allen set-screw was very loose and had screwed itself in.
Ben and I adjusted the switch trip points and verified that the filter
wheel would be able to turn freely. Ben did apply a small amount of
locktite to the set-screw to prevent the failure mode. Bench testing
showed proper operation in and out. Ben cleaned the inside of the unit
with IPA and removed moth and dust debris. We re-installed the unit on
the instrument and tested it using Jon's provided GUI. Detent functions
are working properly.
I have documented the circuitry (attached) that I saw and it is saved
on my computer under my 1m path."
- 3/31/11: Bill replaced 656/25 in slot 3 with 657/27
- 3/28/11: replaced broken dome encoder coupler and adjusted dome motor, after apparent failure of dome tracking.
Ed: "Just wanted to add a couple of things for the documentation of
this repair. Ben also addressed the probable root cause of the flex
coupling failure. The encoder shaft and motor shaft were misaligned so
Ben modified the encoder mounting bracket to proper alignment. We also
re-terminated the encoder connector with a high pressure housing and
receptacles from Molex. Same form factor but the new receptacles grab
the encoder posts from all four sides instead of the previously installed
economy type which only spring tensions along one side."
Ben: "Mark spotted the flexure was cracked in half for the encoder. Ed
found another flexure in your tool box i installed. Also we found the
motor assembly was loose, we tightened it back up. when testing everything
i found that the up limit on the lower hatch was out of adjustment. I
readjusted it and is working correctly."
- 2/24/11: repair of azimuth home sensor.
Ed: "This email simply documents what Ben and I did to the telescope
this morning. We disconnected the TCS rack and brought it into the
lab for servicing. Visual inspection shows clogged fan filters but
the inside of the cabinet looked relatively clean and well laid out.
Re-seat all connectors in the 'X-axis home' signal chain. replace the
IM103 digital output Opto-22 module (red color). Tighten all Opto-22
module retaining screws and blow out dust from cabinet and fan filters.
Reassemble the cabinet and reinstall. Called you up to perform a homing
verification and finished.
Since the Az home problem is very intermittent, only time will tell
if replacing the Opto-22 module repairs the problem. To date we've
re-terminated the magnesensor read head connector (both ends), aligned
the read head and magnet, and replace the Opto-22 module. If it fails
again we'll try replacing the Sony detector (PD-100) or perhaps the
entire kit (read head, magnet, PD-100). I have a request for quote into
Industeq. Inc. for both parts. I'll contact you when the pricing and
availability is known.
Please let me know if the problem returns."
- 10/11/10: Replaced camera after pumping.
Bill: "FYI, vacuum with heat blanket on got down to 5.7e-6
torr. After heat was removed it pumped down to
- late 2007, Mark Klaene and Dave Woods redid the lower dome slit
system to use more robust switches (and relays?).
- 11/06: lube of azimuth bearing
- 05/04: lube of azimuth bearing:
"about 1/3 tube of grease was applied (Mobil fairly lightweight, type ???; this
was done through a small grease port that did not have a grease fitting; we
used such a fitting but the thread of the hole apparently didn't match the
fitting well, so we could not get the fitting to stay in. After lubrication,
a new felt seal was installed using masking tape."